# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ VII-A Error Bound of KernelSHAP
By solving Eq.(A22), we have,
and
Proposition A1.
(The upper bound of $||\bm{\Sigma}^{*{-1}}||$) The $l_{2}$ norm of $\bm{\Sigma}^{*^{-1}}$ is upper bounded by,
$\small||\bm{\Sigma}^{*^{-1}}||\leq\sqrt{\frac{|D|+1}{(\frac{1}{2}-\alpha)^{2}}+\frac{(|D|+1)|D|\alpha^{2}}{((\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2}))^{2}}}.$ (A25)
$\small||\bm{\Sigma}^{*^{-1}}||^{2}\leq||\bm{\Sigma}^{*^{-1}}||^{2}_{F}=\frac{|D|+1}{(\frac{1}{2}-\alpha)^{2}}+\frac{(|D|+1)|D|\alpha^{2}}{((\alpha-\frac{1}{2})(K\alpha-\alpha+\frac{1}{2}))^{2}}.$ (A26)
By taking root square on both sides of Eq.(A26), we can get Eq.(A25).
The upper bound of $||\bm{\Gamma}^{*}||$ can be obtained in a same way.
Proposition A2.
(Upper bound of $||\bm{\Gamma}^{*}||$) The $l_{2}$ norm of $\bm{\Gamma}^{*}$ is upper bounded by,
$\small||\bm{\Gamma}^{*}||\leq\sqrt{|D|+1}.$ (A27)
Proof. The square $l_{2}$ norm of $\bm{\Gamma}^{*}$ is as follows.
$\small||\bm{\Gamma}^{*}||^{2}=||\mathbf{B}\bm{\psi}\mathbf{v}||^{2}.$ (A28)
Since each element in $\bm{\Gamma}^{*}$ is bounded within $[0,1]$, the square $l_{2}$ norm of $\bm{\Gamma}^{*}$ is smaller then its dimension, namely, $||\bm{\Gamma}^{*}||^{2}\leq|D|+1$.
Proposition A3.
(Upper bound of $\hat{\bm{\Sigma}}^{-1}$) For all $\epsilon_{3}\in(0,\frac{1}{||\bm{\Sigma}||_{F}^{*^{-1}}})$, with the probability greater than $1-2(|D|+1)^{2}\exp(-2K\epsilon_{3}^{2})$, there exist a small positive number $\epsilon_{3}^{\prime}$ that
$\small||\hat{\bm{\Sigma}}^{-1}||\leq\frac{||\bm{\Sigma}^{*^{-1}}||_{F}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}=||\bm{\Sigma}^{*^{-1}}||_{F}+\epsilon_{3}^{\prime},$ (A29)
where $\epsilon_{3}^{\prime}=\frac{\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}$.
Proof. According to Proposition A1, we have $\lambda_{\max}(\bm{\Sigma}^{*^{-1}})\leq||\bm{\Sigma}^{*^{-1}}||_{F}$ and
$\small\lambda_{\min}(\bm{\Sigma}^{*})\geq\frac{1}{||\bm{\Sigma}^{*^{-1}}||_{F}}.$ (A30)
According to Corollary  A1, $\hat{\bm{\Sigma}}$ is concentrated on $\bm{\Sigma}^{*}$. By using Weyl inequality [58],
$\small|\lambda_{\min}(\hat{\bm{\Sigma}})-\lambda_{\min}(\bm{\Sigma}^{*})|\leq||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\leq\epsilon_{3},$ (A31)
so that,
$\small\lambda_{\min}(\hat{\bm{\Sigma}})\geq\lambda_{\min}(\bm{\Sigma}^{*})-\epsilon_{3}\geq\frac{1}{||\bm{\Sigma}^{*^{-1}}||_{F}}-\epsilon_{3},$ (A32)
which leads to
$\small||\hat{\bm{\Sigma}}^{-1}||\leq\frac{||\bm{\Sigma}^{*^{-1}}||_{F}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}.$ (A33)
The right hand side of Eq.(A33) can bo obtained as follows.
$\quad\frac{||\bm{\Sigma}^{*^{-1}}||_{F}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}=\frac{||\bm{\Sigma}^{*^{-1}}||_{F}-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}+\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}$ (A34)
$=||\bm{\Sigma}^{*^{-1}}||_{F}+\frac{\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}.$
By defining $\epsilon_{3}^{\prime}=\frac{\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}^{2}}{1-\epsilon_{3}||\bm{\Sigma}^{*^{-1}}||_{F}}$, we can get the upper bound of $||\hat{\bm{\Sigma}}^{-1}||\leq|\hat{\bm{\Sigma}}^{-1}||_{F}+\epsilon_{3}^{\prime}$, with $\epsilon_{3}^{\prime}$ being a small number.
Proof. (Proof of Theorem 2) For fixed feature dimension $|D|$, a small positive $\epsilon_{3}$, and a large sampling number $K$, $\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]$ is upper bounded with the probability greater than $1-2((|D|+1)^{2}+(|D|+1))\exp(-2K\epsilon_{3}^{2})$ as follows.
$\quad\mathbb{E}\left[|\hat{v}(S)-v(S)|\right]\leq\mathbb{E}\left[||\mathbf{b}^{T}||\cdot||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]=(|S|+1)\mathbb{E}\left[||\hat{\bm{\beta}}-\bm{\beta}^{*}||\right]$ (A35)
$\leq\{\mathbb{E}\left[||\hat{\bm{\Sigma}}^{-1}||\cdot||\hat{\bm{\Gamma}}-\bm{\Gamma}^{*}||\right]+\mathbb{E}\left[||\bm{\Sigma}^{*^{-1}}||\cdot||\hat{\bm{\Sigma}}-\bm{\Sigma}^{*}||\cdot||\hat{\bm{\Sigma}}^{-1}||\cdot||\bm{\Gamma}^{*}||\right]$
$\quad+\epsilon_{3}^{2}\}(|S|+1)$
$\leq\{||\bm{\Sigma}^{*^{-1}}||_{F}\cdot\int_{0}^{+\infty}2(|D|+1)\exp(-2K\epsilon_{3}^{2})d\epsilon_{3}+\epsilon_{3}^{2}$
@@
$\cdot(||\bm{\Sigma}^{*^{-1}}||_{F}+\epsilon_{3}^{\prime})\cdot\sqrt{|D|+1}\}(|S|+1)$
$=2(|D|+1)(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||_{F}$
$+2(|D|+1)^{\frac{5}{2}}(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||_{F}(||\bm{\Sigma}^{*^{-1}}||_{F}+\epsilon_{3}^{\prime})+\epsilon_{3}^{2}.$
It can be seen from Eq.(A35) that as the data dimension $|D|$ increases, the error bound of KernelSHAP will increase. On the other hand, by increasing the number of sampling $K$ the error bound will get small.
VII-B Error Bound of VAEAC
Variational autoencoder (VAE) is a typical generative model, which uses a latent variable model to generate a probabilistic representation of the unknown data distribution $p(\mathbf{x})$. The basic idea of the variational autoencoder is to generate the observed variable $\mathbf{x}$ through a Gaussian distributed latent variable $\mathbf{z}$. Specifically, VAE models the joint distribution of $\mathbf{x}$ and $\mathbf{z}$ as follows.
$\small p(\mathbf{x},\mathbf{z})=p(\mathbf{z})p(\mathbf{x}|\mathbf{z}),$ (B1)
Here the conditional distribution $p(\mathbf{x}|\mathbf{z})$ is called an encoder, $\mathbf{z}$ is the latent variable. By marginalizing the latent variable, the data distribution can be approximated.
@@ VII-B Error Bound of VAEAC
This latent variable model helps to get $p(\mathbf{x})$ easily. However, the calculation of the posterior distribution $p(\mathbf{z}|\mathbf{x})$ is always intractable. To address this challenge, VAE constructs a distribution $q(\mathbf{z}|\mathbf{x})$ to approximate $p(\mathbf{z}|\mathbf{x})$. The data log-likelihood of VAE can be expressed as follows.

### v4

@@ VII-B Error Bound of Variational autoencoder
$\qquad\left.\left.+\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{*}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}^{*}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right].$
The first term in the last inequality of Eq.(B8) is called statistical error, and the second is called approximation error. The upper bound of the statistical error is as follows.
$\small\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{*})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right]\leq\frac{\sqrt{\pi}}{\sqrt{2K}}.$ (B9)
The approximation error is related to the amortization gap and the approximation gap, the error introduced from amortization gap can be described as follows.
$\qquad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right]$ (B10)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})d\mathbf{x}_{\bar{S}}-\int f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})d\mathbf{x}_{\bar{S}}\right|\right]$
@@
$\left.\left.-\int f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}-f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}-f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})\right|\right]\right]$
$=\epsilon_{1}.$
The the error introduced from approximation gap can be described as follows,
$\qquad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\right]\right|\right]$
@@
$\qquad\left.\left.-\int f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})d\mathbf{x}_{\bar{S}}\right|\right]$
[…]
$\qquad-\left.\left.\int f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}-f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{*})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{*})}-f(\mathbf{x}_{\bar{S}}^{*},\mathbf{x}_{S}^{*})\right|\right]\right]$ (B11)
$=\delta.$
Thus, the approximation error is bounded with $\epsilon_{1}+\delta$. Finally, the estimation error bound of VAE-based approach is as follows.
$\small\mathbb{E}[|\hat{v}(S)-v(S)|]=\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{1}+\delta.$ (B12)
It can be seen from Eq.(B12), the error bound of VAE-based method is limited by the variational inference. Once $D_{KL}\left(q_{\tau}(\mathbf{z}|\mathbf{x})||p_{\theta}(\mathbf{z}|\mathbf{x})\right)$ becomes large, the error of this method can not be ignored.
VII-C Variance analysis of importance sampling in EmSHAP
In this subsection, we analyze the variance of importance sampling used to estimate the partition function in EmSHAP. The difference between the variance of samples drawn from the proposal conditional distribution and those from the true conditional distribution admits a closed-form expression. Specifically, for samples from the true conditional distribution, the variance is as follows.
$\small\text{Var}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]=\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}^{2}]-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]\right)^{2}.$ (C1)
Meanwhile, for samples drawn from the proposal conditional distribution, the variance of the importance-weighted estimator is,
$\qquad\text{Var}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]$ (C2)
$\qquad-\left(\mathbb{E}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\left[Z_{{\bar{S}};\mathbf{x}_{S}}\frac{p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}{q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}\right]\right)^{2}$
Subtracting Eq.(C2) from Eq.(C1) yields
$\qquad\text{Var}_{\mathbf{x}_{\bar{S}}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]-\text{Var}_{\mathbf{x}_{\bar{S}}\sim q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})}[Z_{{\bar{S}};\mathbf{x}_{S}}]$ (C3)
As can be seen from Eq.(C3), when the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ approaches the true distribution $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, the variance discrepancy between the approximate partition function $Z_{{\bar{S}};\mathbf{x}_{S}}$ and the true partition function diminishes and eventually vanishes.
VII-D Structure of predictive/classification models used in Section VI
In this subsection, the structures of predictive/classification models used in Section VI are presented, which are shown in Tables C1 to C4. All four models are trained using the Adam optimizer with a learning rate of 0.001. The training epoch limit is set to 100 and the training process is stopped when the validation accuracy does not improve for 10 consecutive epochs.
TABLE C1: The structure of CNN for classification of MNIST dataset
Layer Configuration
input $28\times 28\times 1$ grayscale image
conv1 conv1,10,3×3+ReLU+MaxPool
conv2 conv10,20,5×5+ReLU+MaxPool
fc1 fc320,10+Softmax
TABLE C2: The structure of MLP for classification of ADTI dataset
Layer Configuration
input 1212
fc1 fc64,64+ReLU
fc2 fc64,64+ReLU
fc3 fc64,2+Softmax
TABLE C3: The structure of Informer for prediction of ETT dataset
Layer Configuration
input $120\times 6$
encoder ProbSparse self-attention120×512
decoder generative decoder48×512
fc3 fc512,1
TABLE C4: The structure of CAML for classification of MIMIC dataset
Layer Configuration
input 2500
embedding $10947\times 100$
dropout dropout(0.5)
conv1 conv100,100,5×5,padding=2+tanh
attention fc100,50+softmax+dot-product
fc1 fc100,50
VII-E More details for feature attribution in MNIST and MIMIC-III datasets
In this subsection, additional experimental results using different feature attribution methods on the MNIST and MIMIC-III datasets are presented.
VII-E1 Feature attribution on MNIST dataset under noisy environment
To test the performance of the proposed method under noisy data, Gaussian distributed noises with zero mean and varying degrees of variance are added to the original images. The variances vary from 0.1 to 0.5, with an interval of 0.1. Feature attribution is then performed on these noisy images. Figs.D1 and D2 present the attribution results produced by the six methods on the noisy images.
Figure D1: Feature attributions for noise-polluted image of digit 3. From top to bottom: noise variances of 0.1, 0.2, 0.3, 0.4, and 0.5.
Figure D2: Feature attributions for noise-polluted image of digit 8. From top to bottom: noise variances of 0.1, 0.2, 0.3, 0.4, and 0.5.
It can be seen from Figs. D1 and D2 that when the variances of noise increase, the feature attribution performance of all methods deteriorates. Whilst the sharpest performance deterioration can be observed for SamplingSHAP, the performance for KernelSHAP, ST-SHAP, and EmSHAP are relatively stable. This verifies the robustness of EmSHAP in the case of noise.
VII-E2 Feature attribution on MIMIC-III dataset
This subsection presents additional feature attribution results on the MIMIC-III dataset to further validate the effectiveness of EmSHAP. Figs. D3 and D4 present the attribution results on two additional cases from the MIMIC-III dataset, corresponding to subject IDs 102024 and 136168, respectively. For subject ID 102024, the predicted codes include ”end stage renal disease” and ”hyperpotassemia”. EmSHAP correctly highlights clinically meaningful terms such as ”lupus”, ”hyperkalemia”, and ”Kayexalate”, which are directly associated with the patient’s renal dysfunction and electrolyte imbalance. Similarly, for subject ID 136168, the model predicts ”coronary atherosclerosis of native coronary artery” and ”paroxysmal ventricular tachycardia”. EmSHAP assigns high attribution scores to terms such as ”myocardial, ”ST elevation”, ”electrocardiogram”, and ”right atrial”, which are well recognized as indicative of acute myocardial infarction and arrhythmia, thereby demonstrating consistency with medical knowledge.
Figure D3: Feature attribution results on MIMIC-III dataset (subject ID: 102024).
Figure D4: Feature attribution results on MIMIC-III dataset (subject ID: 136168).
Despite the better feature attribution results, EmSHAP also spuriously produces attributions to general high-frequency words with limited clinical meaning (e.g., general descriptors like ”she” or ”an”). This likely arises due to the model’s sensitivity to frequent background terms in clinical narratives. While these cases indicate that the method may not always yield perfectly precise explanations, EmSHAP nevertheless is proved to be capable of capturing clinically important signals in large-scale, high-dimensional datasets such as MIMIC-III. These results suggest that EmSHAP offers practically useful interpretability for automated medical coding by identifying the key textual features driving predictions under complex real-world conditions.
VII-F Performance analysis of EmSHAP under different parameters and model structures
TABLE E1: SIC AUC results for different masking rates
Dataset
SIC
AUC
Masking rate
0.5
0.6
0.7
0.8
0.9
Dynamic
masking
MNIST